OpenStudy (anonymous):

factor completely 10x^3 + 80

OpenStudy (anonymous):

first is 10(x^3+8)

OpenStudy (anonymous):

you should know the rest

OpenStudy (zzr0ck3r):

hint: 2^3 = 8

OpenStudy (zzr0ck3r):

a^3 + b^3 = (a + b)^3 - 3 a b (a + b)

OpenStudy (anonymous):

(x+2) (x^2 -2x +4)?

OpenStudy (anonymous):

10* before (x + 2)

OpenStudy (anonymous):

\[10 x^3+80=10 (x+2) \left(x^2-2 x+4\right) \]

OpenStudy (anonymous):

@Jenny4218 is correct

OpenStudy (anonymous):

tank you